A 50-N force acts on a 2-kg crate that starts from rest. There is no friction. At time equals 2.0 seconds, the instantaneous rate at which the force is doing work is:

Grade:12th pass

3 Answers

Arun
3 years ago
Dear student
 
Thereis no image attached. Please repost the question with an attachment. I will be happy to help you.
Mayank
3 years ago
I am sorry sir I do not have image for this question and if u can manage without image would be greately appreciated
 
Vikas TU
3 years ago
Dear student 
instantaneous Power delivered by a moving force equals:
P = force x velocity
The final velocity in this case is
Vf = (1/m)*(final momentum) = F*t/m
= 50*2/2 = 50 m/s
Power at end of interval = F*Vf = 50*50 = 2500 W